OpenStudy (anonymous):

find the area of the triangle with a base of 1 2/3 yd and height 3/7 yd

OpenStudy (anonymous):

convert fractions to improper so base = 5/3 yards and height = 3/7 yards, now use the equation Area=(1/2)(Base)(Height) (all multiplied)

OpenStudy (anonymous):

you should get 5/14 yards^2